# Breaking the Gaussian Barrier: Residual-PAC Privacy for Automatic Privatization

### v1

@@ Appendix A Proof of Proposition 2
Setting $\frac{\partial\mathcal{L}}{\partial\ell_{i}}=0$ gives $1=\lambda\frac{r_{i}}{2\ell_{i}(\ell_{i}+r_{i})}$Rightarrow $2\ell_{i}(\ell_{i},r_{i})=\lambda r_{i}$. Equivalently, $\ell^{2}_{i}+r_{i}\ell_{i}-\lambda\frac{r_{i}}{2}=0$, which gives a unique $\ell_{i}(\lambda)=\frac{-r_{i}+\sqrt{r^{2}_{i}+2\lambda r_{i}}}{2}>0$.
@@
Let $F(\lambda)=\frac{1}{2}\sum^{d}_{i=1}\log(1+\frac{r_{i}}{\ell_{i}(\lambda)})$. We can check that as $\lambda\rightarrow 0^{+}$, each $\ell_{i}(\lambda)\rightarrow 0^{+}$, so $F(\lambda)\rightarrow+\infty$. As $\lambda\rightarrow+\infty$, each $\ell_{i}(\lambda)\rightarrow+\infty$, leading to $F(\lambda)\rightarrow+\infty$. In addition, $\frac{dF(\lambda)}{d\lambda}<0$ throughout, so $F$ is strictly decreasing from +$\infty$+ ∞ down to 00. Therefore, there is a unique $\lambda^{*}>0$ such that $F(\lambda^{*})=\beta$. At this $\lambda^{*}$, each $\ell^{*}_{i}=\ell^{*}_{i}(\lambda^{*})$ is unique. Thus, $\Sigma^{*}_{B}=\hat{U}\mathrm{diag}(\ell^{*}_{1},\dots,\ell^{*}_{d})\hat{U}^{T}$ is unique minimizer of $\mathrm{tr}(\Sigma_{B})$. By construction, $\frac{1}{2}\log\mathrm{det}\left(1+\Sigma_{\mathcal{M}(X)}(\Sigma^{*}_{B})^{-1}\right)=\beta$. Therefore, it is also the unique minimizer of (5). ∎
Appendix B Proof of Theorem 3
@@
Appendix C Proof of Theorem 4
We first prove the inequality chain, and then prove the strict positivity $\widehat{\mathtt{D}}_{Z}>0$.
C.1 Part 1: Inequality Chain
Define the Stein score
$T(\tilde{Z})=\nabla_{\tilde{z}}\log P_{\widetilde{Z}}(\tilde{z})+\Sigma^{-1}_{\widetilde{Z}}\tilde{z}.$
For any distribution $P$ on $\mathbb{R}^{d}$ with density $p$, mean $\mu$, and covariance $\Sigma$, and the moment-matched Gaussian $g=\mathcal{N}(\mu,\Sigma)$, the Gaussian log-Sobolev inequality gives
$\mathtt{D}_{\mathrm{KL}}(P\|g)\geq\frac{1}{2}\mathbb{E}_{P}\left[\|T(Z_{p})\|^{2}_{2}\right]\equiv\frac{1}{2}\mathcal{J}(P\|g),$ (10)
where $Z_{p}\sim P$.
Let $s_{\alpha^{*}}$ be the DSM-optimal score estimator trained on the centered samples $\{\tilde{z}_{i}\}^{N}_{i=1}$. By [38] and [39], $s_{\alpha^{*}}(\tilde{z})$ converges to $\Delta_{\tilde{z}}\log P_{\widetilde{Z}}(\tilde{z})$ in $L^{2}(P_{\widetilde{Z}})$ as $N\rightarrow\infty$ and $\epsilon\rightarrow 0$.
Taking outer expectation over the i.i.d. draws of the sample $\{\tilde{z}_{i}\}$ and the DSM training randomness yields
$\mathbb{E}[\widehat{\mathtt{D}}_{Z}]=\frac{1}{2}\mathbb{E}_{\widehat{Z}]}\left[\|s_{\alpha^{*}}(\widehat{Z})\|^{2}_{2}\right]-\mathbb{E}_{\widehat{Z}}\left[\nabla\cdot s_{\alpha^{*}}(\widetilde{Z})\right].$
For any square0integrable vector filds $f$, by the Stein’ identity, we have
$\mathbb{E}_{\widetilde{Z}}\left[\nabla\cdot f(\widetilde{Z})\right]=-\mathbb{E}_{\widetilde{Z}}\left[f(\widetilde{Z})^{\top}\Sigma^{-1}_{\tilde{Z}}\widetilde{Z}\right].$
When $f=s_{\alpha^{*}}$, we get
$\mathbb{E}[\widehat{\mathtt{D}}_{Z}]=\frac{1}{2}\mathbb{E}_{\widehat{Z}]}\left[\|s_{\alpha^{*}}(\widehat{Z})\|^{2}_{2}+2s^{\top}_{\alpha^{*}}\Sigma^{-1}_{\widetilde{Z}}\widetilde{Z}\right].$
Adding and subtracting $\frac{1}{2}\|\Sigma^{-1/2}_{\widetilde{Z}}\widetilde{Z}\|^{2}_{2}$ gives
$\mathbb{E}\left[\widehat{\mathtt{D}}_{Z}\right]=\frac{1}{2}\mathbb{E}\left[\|s_{\alpha^{*}}(\widetilde{Z})+\Sigma^{-1}_{\widetilde{Z}}\widetilde{Z}\|^{2}_{2}\right]-\frac{1}{2}\mathbb{E}\left[\|\Sigma^{-1/2}_{\widetilde{Z}}\widetilde{Z}\|^{2}_{2}\right].$
Since $\mu_{\tilde{Z}}=0$ and $\Sigma_{\widetilde{Z}}=\Sigma_{Z}\succ 0$,
$\frac{1}{2}\mathbb{E}\left[\|\Sigma^{-1/2}_{\widetilde{Z}}\widetilde{Z}\|^{2}_{2}\right]=\frac{d}{2}$
is a constant that is independent of $s_{\alpha^{*}}$. Thus, we have
$\mathbb{E}\left[\widehat{\mathtt{D}}_{Z}\right]\leq\frac{1}{2}\mathbb{E}\left[\|s_{\alpha^{*}}(\widetilde{Z})+\Sigma^{-1}_{\widetilde{Z}}\widetilde{Z}\|^{2}_{2}\right]=\frac{1}{2}\mathbb{E}\left[\|T_{\alpha^{*}}(\widetilde{Z})\|^{2}_{2}\right],$
where $T_{\alpha^{*}}(\tilde{z})\equiv s_{\alpha^{*}}(\tilde{z})+\Sigma^{-1}_{\widetilde{Z}}\tilde{z}$.
Therefore, as $N\rightarrow\infty$ and $\epsilon\rightarrow 0$, we have
$\lim_{N\rightarrow\infty,\epsilon\rightarrow 0}\mathbb{E}\left[\widehat{\mathtt{D}}_{Z}\right]\leq\frac{1}{2}\mathbb{E}\left[\|T(\widetilde{Z})\|^{2}_{2}\right]=\frac{1}{2}\mathcal{J}(P_{\widetilde{Z}}\|\mathcal{N}).$
By (10), $\frac{1}{2}\mathcal{J}(P_{\widetilde{Z}}\|\mathcal{N})\leq\mathtt{D}_{Z}$. Therefore, we obtain the inequality chain
$\mathbb{E}\left[\widehat{\mathtt{D}}_{Z}\right]\leq\frac{1}{2}\mathbb{E}\left[\|\nabla_{z}\ln p_{Z}(Z)\|^{2}_{2}\right]\leq\mathtt{D}_{Z}.$
C.2 Part 2: Strict Positivity
For $\epsilon>0$, define
$\tilde{s}(\tilde{z})=s_{\alpha^{*}}(\widetilde{z})+(\widehat{\Sigma}_{Z}^{-1}+\epsilon I)\widetilde{z}$
Then, $T_{\tilde{s}}(\tilde{z})=\tilde{s}(\tilde{z})+\widehat{\Sigma}^{-1}_{Z}\tilde{z}$ differs from zero by at least the term $\epsilon\tilde{z}$. Unless in the degenerate case when $\widehat{Z}$ is almost surely zero,
$\mathbb{E}[\|T_{\tilde{s}}(\widetilde{Z})\|^{2}_{2}]\geq\epsilon^{2}\mathbb{E}[\|\widetilde{Z}\|^{2}_{2}]>0.$
Here, $\mathbb{E}[\|T_{\tilde{s}}(\widetilde{Z})\|^{2}_{2}]\geq d$, where the equality holds only for Gaussian $\widetilde{Z}$. Then, from
$\widehat{\mathtt{D}}_{Z}=\frac{1}{2}\|T_{\tilde{s}}(\tilde{z})\|^{2}_{2}-\frac{d}{2},$
we guarantee $\widehat{\mathtt{D}}_{Z}>0$ for non-Gaussian $Z$.
Appendix D Proof of Theorem 5
For any real-valued random variable $Z$ with mean $\mu_{z}$ and variance $\sigma^{2}_{Z}<\infty$, write the excess kurtosis,
$\kappa_{4}=\mathbb{E}[(Z-\mu_{Z})^{2}]-3\sigma^{4}_{Z}.$
By the fourth-moment bund [44], the KL divergence between $P_{Z}$ and the moment-matched Gaussian satisfies
$\mathtt{D}_{\mathrm{KL}}(P_{Z}\|\mathcal{N}(\mu_{Z},\sigma^{2}_{Z}))\geq\frac{\kappa^{2}}{48\sigma^{4}_{Z}}.$ (11)
With i.i.d. samples $\{z_{i}\}=1$, define the centered samples $\tilde{z}_{i}=z_{i}-\mu_{z}$ and the empirical moments
$\tilde{\sigma}_{Z}=\frac{1}{N}\sum^{N}_{i=1}\tilde{z}^{2}_{i},\;\tilde{\kappa}_{4}=\frac{1}{N}\sum^{N}_{i=1}\tilde{z}^{2}_{i}-3(\tilde{\sigma}^{2}_{Z})^{2}.$
By the strong law of large numbers, we have
[…]
For a fixed $c>0$, let $\widehat{\mathtt{D}}_{Z}=\frac{\left(\max\left(\widehat{\kappa}_{4},\frac{c}{N}\right)\right)^{2}}{48(\widehat{\sigma}_{Z}^{2})^{2}}$, where the numerator is the square of a non-negative quantity and the denominator is a positive empirical variance. Thus, $\widehat{\mathtt{D}}_{Z}>0$ almost surely for every finite $N$.
[…]
Appendix E Proof of Proposition 4
@@
[…]
$=\int_{\mathcal{X},\mathcal{Y},\mathcal{R}^{d}}P_{X}(x)G_{\mathcal{M},Q}(y|x,b)\log\frac{\eta_{Q}(x|y+b)}{\pi(x|y+b)}dxdydb.$
@@
[…]
@@
Appendix F Proof of Proposition 5
@@
Suppose, to reach a contradiction, that an optimal $Q^{\star}$ is isotropic with $\Sigma_{Q^{\star}}=\sigma^{2}I_{d}$ and attains the constraint with equality: $\mathcal{H}(X\mid\mathcal{M}(X)+B)=\hat{\beta}$.
@@
Therefore $\mathcal{K}(Q^{\prime})-\mathcal{K}(Q^{\star})<g^{\prime}(\sigma^{2})(\Delta_{u}-\Delta_{v})<0.$ That is, $Q^{\prime}$ is feasible and cheaper than $Q^{\star}$, contradicting optimality. Hence no optimum can be isotropic, so every minimiser must have $\lambda_{\max}(\Sigma)>\lambda_{\min}(\Sigma)$.
Appendix G Proof of Proposition 6
G.1 Part (i):
@@ G.2 Part (ii):
$1+\lambda g(w)=0\quad\text{if }\sigma_{w}^{2}>0,\;1+\lambda g(w)\geq 0\quad\text{if }\sigma_{w}^{2}=0,$
@@
these equalities can hold only for as long as the required mutual information reduction does not exceed $\beta_{\mathtt{lab}}$. Therefore, $\sigma_{v}^{2}=0$ for every $v\in S_{\mathtt{lab}}$. With those label‑directions undisturbed, each class margin $e_{\ell}-e_{j}$ retains its sign, whence $\arg\max_{i}(Z_{i}+B^{\star}_{i})=\hat{y}$.
Appendix H Proof of Theorem 6
H.1 Part (i)
@@
H.2 Part (ii)
It is well known that for a fixed prior, mutual information is convex in the channel law. When $Z=\mathcal{M}(X)+B$, the “channel law” in our setting of the deterministic mechanism is determined by the perturbation rule $Q$. Thus, the mapping $Q\mapsto\mathtt{MI}(Q)\equiv\mathtt{MI}(X;\mathcal{M}(X)+B$ is convex. The objective $\mathcal{K}(Q)=\mathbb{E}_{Q}[\|B\|^{2}_{2}]$ is linear (hence convex) in $Q$. In addition, the constraint set $\{Q:\mathtt{MI}(Q)\leq\beta\}$ is convex. Then, Slater’s condition holds because:

### v2

@@ Appendix I Proof of Corollary 2
Setting $\frac{\partial\mathcal{L}}{\partial\ell_{i}}=0$ gives
$1=\lambda\frac{r_{i}}{2\ell_{i}(\ell_{i}+r_{i})}\Rightarrow 2\ell_{i}(\ell_{i},r_{i})=\lambda r_{i}.$
Equivalently, $\ell^{2}_{i}+r_{i}\ell_{i}-\lambda\frac{r_{i}}{2}=0$, which gives a unique $\ell_{i}(\lambda)=\frac{-r_{i}+\sqrt{r^{2}_{i}+2\lambda r_{i}}}{2}>0$.
Let
$F(\lambda)=\frac{1}{2}\sum^{d}_{i=1}\log(1+\frac{r_{i}}{\ell_{i}(\lambda)}).$
We can have the following:
  • As $\lambda\rightarrow 0^{+}$, each $\ell_{i}(\lambda)\rightarrow 0^{+}$, leading to $F(\lambda)\rightarrow+\infty$.
  • As $\lambda\rightarrow+\infty$, each $\ell_{i}(\lambda)\rightarrow+\infty$, leading to $F(\lambda)\rightarrow+\infty$.
@@
In addition, $\frac{dF(\lambda)}{d\lambda}<0$ throughout. Thus, $F$ is strictly decreasing from $\infty$ down to 0. Therefore, there is a unique $\lambda^{*}>0$ such that $F(\lambda^{*})=\beta$. At this $\lambda^{*}$, each $\ell^{*}_{i}=\ell^{*}_{i}(\lambda^{*})$ is unique. Thus, $\Sigma^{*}_{B}=\hat{U}\mathrm{diag}(\ell^{*}_{1},\dots,\ell^{*}_{d})\hat{U}^{T}$ is unique minimizer of $\mathrm{tr}(\Sigma_{B})$. By construction,
$\frac{1}{2}\log\mathrm{det}\left(1+\Sigma_{\mathcal{M}(X)}(\Sigma^{*}_{B})^{-1}\right)=\beta.$
Therefore, it is also the unique minimizer of (5). ∎
Appendix L Proof of Proposition 3
For additive Gaussian noise with covariance $\Sigma_{B}\succ 0$,
$\mathtt{MI}(X;\mathcal{M}(X)+B)\ \leq\ \tfrac{1}{2}\,\log\det\big{(}I+\Sigma_{M}\Sigma_{B}^{-1}\big{)}.$
The trace–optimal (eigen-aligned) choice that enforces the log-det constraint at level $\hat{\beta}$ has eigenvalues $e_{i}^{*}=\alpha\sqrt{\lambda_{i}}$ in the UU–basis, with $\alpha=\frac{S}{2\hat{\beta}}$. Hence
$\mathbb{E}\|B\|_{2}^{2}=\operatorname{tr}(\Sigma_{B}^{*})=\sum_{i}e_{i}^{*}=\frac{S^{2}}{2\hat{\beta}}$.
Algorithm 1 constructs a diagonal precision $\Lambda_{B}$ in the empirical eigenbasis and, when the "gap" test passes, let
$\lambda_{B,i}=\frac{2v}{\sqrt{\hat{\lambda}_{i}+\delta}\ \sum_{k}\sqrt{\hat{\lambda}_{k}+\delta}}\text{ and }\delta=\frac{10cv}{\beta}.$
Plugging the population spectrum ($\hat{\lambda}=\lambda$, same $U$) yields noise eigenvalues
$e_{i}^{\gamma}=\frac{\sqrt{\lambda_{i}+\delta}\ \sum_{k}\sqrt{\lambda_{k}+\delta}}{2v},$
and
$\operatorname{tr}(\Sigma_{B}^{\gamma})=\frac{\Big{(}\sum_{k}\sqrt{\lambda_{k}+\delta}\Big{)}^{2}}{2v}.$
Since $\delta\geq 0$ and $v\leq\hat{\beta}$, we have
$\operatorname{tr}(\Sigma_{B}^{\gamma})\geq\frac{S^{2}}{2v}\geq\frac{S^{2}}{2\hat{\beta}}=\operatorname{tr}(\Sigma_{B}^{*}),$
which proves (i) in this branch by (L)–(L). Moreover,
$\frac{e_{i}^{\gamma}}{e_{i}^{*}}=\frac{\sum_{k}\sqrt{\lambda_{k}+\delta}}{\sum_{k}\sqrt{\lambda_{k}}}\frac{\hat{\beta}}{v}\sqrt{1+\frac{\delta}{\lambda_{i}}}\ \geq\ 1.$
Thus, $\Sigma_{B}^{\gamma}\succeq\Sigma_{B}^{*}$. For additive Gaussian noise perturbation, increasing $\Sigma_{B}$ (in positive semidefinite order) implies the log-det bound decreases, hence the mutual information decreases:
$\mathtt{MI}\big{(}X;\mathcal{M}(X)+B_{\gamma}\big{)}$ $\leq\tfrac{1}{2}\log\det\big{(}I+\Sigma_{M}(\Sigma_{B}^{\gamma})^{-1}\big{)}$
$\leq\tfrac{1}{2}\log\det\big{(}I+\Sigma_{M}(\Sigma_{B}^{*})^{-1}\big{)}.$
Since the Auto-PAC design saturates the bound at $\hat{\beta}$ (and the true MI is bounded by it), we obtain
$\mathtt{MI}\big{(}X;\mathcal{M}(X)+B_{\gamma}\big{)}\ \leq\ \mathtt{MI}\big{(}X;\mathcal{M}(X)+B\big{)},$
establishing (ii) in this branch.
When the "gap" test fails, Algorithm 1 uses $\Sigma_{B}^{\mathrm{iso}}=\alpha I$ with $\alpha=(\sum_{i}\hat{\lambda}_{i}+dc)/(2v)$, hence
$\operatorname{tr}(\Sigma_{B}^{\mathrm{iso}})=\frac{d}{2v}\Big{(}\sum_{i}\lambda_{i}+dc\Big{)}\ \geq\ \frac{1}{2v}\,S^{2}\ \geq\ \frac{S^{2}}{2\hat{\beta}}=\operatorname{tr}(\Sigma_{B}^{*}),$
where we used Cauchy–Schwarz $S^{2}=(\sum_{i}\sqrt{\lambda_{i}})^{2}\leq d\sum_{i}\lambda_{i}$ and $v\leq\hat{\beta}$. Thus (i) also holds in this branch.
For (ii), both designs enforce the same budget $\hat{\beta}$:
$\mathtt{MI}\big{(}X;\mathcal{M}(X)+B_{\gamma}\big{)}\ \leq\ \hat{\beta},\qquad\mathtt{MI}\big{(}X;\mathcal{M}(X)+B\big{)}\ \leq\ \hat{\beta}$.
In the distinct–eigenvalues branch we proved the stronger order $\leq$ between the two MI’s. In the isotropic fallback, the same order holds whenever $\alpha I\succeq\Sigma_{B}^{*}$ (e.g., for nearly isotropic spectra); otherwise we keep the common upper bound $\hat{\beta}$. Either way, the stated inequality (ii) is satisfied in the branch where Algorithm 1’s eigenbasis matches $U$, and the confidence guarantee always preserves the budget. ∎
Appendix M Proof of Proposition 4
Since $\beta_{1}<\beta_{2}$, any distribution $Q$ satisfying $I_{\mathrm{true}}(Q)\leq\beta_{1}$ necessarily satisfies $I_{\mathrm{true}}(Q)\leq\beta_{2}$. Consequently, we have the inclusion $\mathcal{F}(\beta_{1})\subseteq\mathcal{F}(\beta_{2})$. Let $A$ and $\hat{A}$ be arbitrary sets with $A\subseteq\hat{B}$, and let $f$ be any real-valued function defined on $B$. Then, $\inf_{x\in A}f(x)\geq\inf_{x\in\hat{A}}f(x)$, with equality holding when the infimum over $\hat{A}$ is attained within the subset $A$. Applying this with $A=\mathcal{A}(\beta_{1})$, $\hat{A}=\mathcal{F}(\beta_{2})$, and $f(Q)=\mathbb{E}_{Q}\bigl{[}\|B\|_{2}^{2}\bigr{]}$ yields
$\inf_{Q\in\mathcal{F}(\beta_{1})}\mathbb{E}_{Q}\bigl{[}\|B\|_{2}^{2}\bigr{]}\geq\inf_{Q\in\mathcal{F}(\beta_{2})}\mathbb{E}_{Q}\bigl{[}\|B\|_{2}^{2}\bigr{]}$
By definition, $Q^{*}(\beta_{i})$ achieves the infimum of $\mathbb{E}_{Q}\bigl{[}\|B\|_{2}^{2}\bigr{]}$ over $\mathcal{F}(\beta_{i})$ for $i=1,2$. Therefore,
$\mathbb{E}_{Q^{*}(\beta_{1})}\bigl{[}\|B\|_{2}^{2}\bigr{]}$ $=\inf_{Q\in\mathcal{F}(\beta_{1})}\mathbb{E}_{Q}\bigl{[}\|B\|_{2}^{2}\bigr{]}\geq\inf_{Q\in\mathcal{F}(\beta_{2})}\mathbb{E}_{Q}\bigl{[}\|B\|_{2}^{2}\bigr{]}$
$=\mathbb{E}_{Q^{*}(\beta_{2})}\bigl{[}\|B\|_{2}^{2}\bigr{]}.$
Appendix N Proof of Theorem 2
@@
Appendix O Proof of Proposition 5
@@
$=\int_{\mathcal{X},\mathcal{Y},\mathbb{R}^{d}}P_{X}(x)G_{\mathcal{M},Q}(y|x,b)\log\eta_{Q}(x|y+b)dxdydb$
$-\int_{\mathcal{X},\mathcal{Y},\mathbb{R}^{d}}P_{X}(x)G_{\mathcal{M},Q}(y|x,b)\log\pi(x|y+b)dxdydb$
$=\int_{\mathcal{X},\mathcal{Y},\mathbb{R}^{d}}P_{X}(x)G_{\mathcal{M},Q}(y|x,b)\log\frac{\eta_{Q}(x|y+b)}{\pi(x|y+b)}dxdydb.$
@@
$\mathbf{P}_{Q}(y)\equiv\int_{\mathcal{X},\mathbb{R}^{d}}P_{X}(x)G_{\mathcal{M},Q}(y|x,b)dxdb.$
@@
Appendix P Proof of Proposition 6
@@
Suppose, to reach a contradiction, that an optimal $Q^{*}$ is isotropic with $\Sigma_{Q^{*}}=\sigma^{2}I_{d}$ and attains the constraint with equality: $\mathcal{H}(X\mid\mathcal{M}(X)+B)=\hat{\beta}$.
@@
Therefore $\mathcal{K}(Q^{\prime})-\mathcal{K}(Q^{*})<g^{\prime}(\sigma^{2})(\Delta_{u}-\Delta_{v})<0.$ That is, $Q^{\prime}$ is feasible and cheaper than $Q^{*}$, contradicting optimality. Hence no optimum can be isotropic, so every minimiser must have $\lambda_{\max}(\Sigma)>\lambda_{\min}(\Sigma)$.
Appendix Q Proof of Proposition 7
Q.1 Part (i):
@@
[…]
@@
these equalities can hold only for as long as the required mutual information reduction does not exceed $\beta_{\mathtt{lab}}$. Therefore, $\sigma_{v}^{2}=0$ for every $v\in S_{\mathtt{lab}}$. With those label-directions undisturbed, each class margin $e_{\ell}-e_{j}$ retains its sign, whence $\arg\max_{i}(Z_{i}+B^{*}_{i})=\hat{y}$.
Appendix R Proof of Theorem 5
R.1 Part (i)
@@
R.2 Part (ii)
It is well known that for a fixed prior, mutual information is convex in the channel law. When $Z=\mathcal{M}(X)+B$, the “channel law" in our setting of the deterministic mechanism is determined by the perturbation rule $Q$. Thus, the mapping $Q\mapsto\mathtt{MI}(Q)\equiv\mathtt{MI}(X;\mathcal{M}(X)+B$ is convex. The objective $\mathcal{K}(Q)=\mathbb{E}_{Q}[\|B\|^{2}_{2}]$ is linear (hence convex) in $Q$. In addition, the constraint set $\{Q:\mathtt{MI}(Q)\leq\beta\}$ is convex. Then, Slater’s condition holds because: